What mass of steam at 100°C must be added to 1.84 kg of ice at 0°C to yield liquid water at 28°C?
Answer #1

heat supplied by stem = heat gained by ice to convert to water

let the mass of steam = m

L = latent heat of vaporization

m (2,260 kJ/kg ) + m ( 4.19 kj/kg) ( 100-28) = 1.84 ( 334 Kj/kg) + 1.84 ( 28 ) ( 4.19 kJ/kg)

m ( 2561.68) = 1.84 ( 451.32)

m = 0.324 kg apprx